OpenStudy (anonymous):

Use the circle graph shown below to answer the question. If 210 people said football was their favorite sport to watch, how many people were surveyed.

OpenStudy (anonymous):

 42% football 33% baseball 25% soccer

OpenStudy (anonymous):

My graph is a little off, can someone please explain how I would solve this problem?

OpenStudy (mertsj):

42 percent of the total choose football Right?

OpenStudy (anonymous):

Yes!

OpenStudy (mertsj):

210 people choose football right?

OpenStudy (anonymous):

Yes!

OpenStudy (mertsj):

Then 42 percent OF the total = 210 right?

OpenStudy (anonymous):

Right :)

OpenStudy (mertsj):

Do you know that in mathematics, "of" means times?

OpenStudy (anonymous):

multiplication, yes!

OpenStudy (mertsj):

So we have: 42% of total = 210

OpenStudy (mertsj):

or: .42t=210

OpenStudy (mertsj):

Solve that equation where t is the total number of people in the survey

OpenStudy (anonymous):

So divide .42 from both sides to get t= 210/.42 ??

OpenStudy (anonymous):

But @Mertsj that equals .002 that can't be right ...

OpenStudy (mertsj):

\[\frac{210}{.42}\times \frac{100}{100}=\frac{21000}{42}=500\]

OpenStudy (mertsj):

You need to type 210 divided by .42 not the other way around. You divided .42 by 210

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} quantity&percentage\\ \hline\\ 210&42\\ x&100 \end{array}\implies \bf \cfrac{210}{x}=\cfrac{42}{100}\)

OpenStudy (anonymous):

Thank you! Got it! Does the 100 represent 100% on the pie chart?

OpenStudy (mertsj):

yw